OpenStudy (anonymous):

I NEED HELP BAD!!!!!!! Evaluate the expression 3 ¼ + 2 5/8 Show your work please!!!

OpenStudy (anonymous):

\[3\frac{ 1 }{ 4 }+2\frac{ 5 }{ 8 }\]\[3\frac{ 1*2 }{ 4*2 }+2\frac{ 5 }{ 8 }\]\[3\frac{ 2 }{ 8 }+2\frac{ 5 }{ 8 }\]\[5\frac{ 7 }{ 8 }\]

OpenStudy (anonymous):

Thank you thank you so much!!! By the way, your profile pic is soooo cute! :)